Suppose the tube in the previous problem is kept vertical with B upward. Water enters through B at the rate of 1 cm3 s-1. Repeat parts (a), (b) and (c). Note that the speed decrease as the water falls down.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. 1/2 pvB2 + pghB + pB = 1/2 pvA2 + pghA + pA 31. 1/2 pvA2 + pghA + pA =1/2 pvB2 + pghB + pB ⇒PB – PA = 1/2 p(vA2 – vB2) + pg (hA – hB)